OpenStudy (anonymous):

add y/y^2-3y-4 + 5/y+1

OpenStudy (anonymous):

lcm =y^2-3y-4 = (y-4)(y+1) we have y + 5(y-4) / y^2-3y-4 = 6y - 20 / y^2-3y-4